The earth is, on average, 150 million kilometers from the sun. Calculate its average speed in orbit.


1
Expert's answer
2021-10-25T17:54:03-0400
"v^2=G\frac{M}{R}\\v^2=6.67\cdot10^{-11}\frac{2\cdot10^{30}}{1.5\cdot10^{11}}\\v=29.8\frac{km}{s}"